Find the unit tangent T, the unit normal N and the curvature K for the curve r(t) =<2 sint, 5t, 2 cost>.
1
Expert's answer
2020-09-24T17:04:39-0400

"\vec{r}(t) = \lbrace 2\sin t, 5t, 2\cos t \rbrace."

Let us take derivatives with respect to t:

"\vec{r}'(t) = \lbrace 2\cos t, 5, -2\sin t \rbrace, \;\; \vec{r}''(t) = \lbrace -2\sin t, 0, -2\cos t \rbrace."


The tangent is "\vec{\tau}(t) = \vec{r}'(t) = \lbrace 2\cos t, 5, -2\sin t \rbrace." We should normalize this vector (norm is "\sqrt{4\cos^2t + 25 + 4\sin^2t} = \sqrt{29}" ), so the unit tangent is "\vec{T} =\left \lbrace \dfrac{2\cos t}{\sqrt{29}}, \dfrac{5}{\sqrt{29}}, \dfrac{-2\sin t}{\sqrt{29}} \right\rbrace" .


The normal is "\vec{\beta}(t) = \vec{r}'(t)\times (\vec{r}'(t)\times\vec{r}''(t) ) ."

"\vec{r}'(t)\times\vec{r}''(t) = \begin{vmatrix}
\vec{i} & \vec{j} & \vec{k} \\
 2\cos t & 5 & -2\sin t \\
-2\sin t & 0 & -2\cos t
 \end{vmatrix}
 = \lbrace -10\cos t, 4 , 10\sin t \rbrace ."

"\vec{r}'(t)\times (\vec{r}'(t)\times\vec{r}''(t) ) = \begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 2\cos t & 5 & -2\sin t \\
 -10\cos t & 4 & 10\sin t
\end{vmatrix} = \lbrace 58\cos t, 0, 58\sin t \rbrace."

The unit normal is "\vec{N} = \lbrace \cos t, 0, \sin t \rbrace."


The curvature is "k = \dfrac{||\vec{r}'(t)\times\vec{r}''(t)||}{||\vec{r}'(t)||^3} = \dfrac{\sqrt{116}}{\sqrt{29}^3} = \dfrac{2}{29}."